Storage cluster issues
Incident Report for Flying Circus
Resolved
The crashed Ceph OSDs and the crashed VM server have been behaving properly over the last hours.
Posted about 1 month ago. May 16, 2017 - 21:10 CEST
Monitoring
A fix has been implemented and we are monitoring the results.
Posted about 1 month ago. May 16, 2017 - 15:35 CEST
Update
Both the storage cluster and the VM host are up again, most sites are working now. Still investigating...
Posted about 1 month ago. May 16, 2017 - 15:31 CEST
Update
Storage is up again, but there seem to be still issues with a VM host.
Posted about 1 month ago. May 16, 2017 - 15:11 CEST
Investigating
We are currently seeing slow I/O requests on the main storage cluster due to a yet undetermined problem.
Posted about 1 month ago. May 16, 2017 - 14:58 CEST